Find the equation in slope-intercept form that is parallel to y=x-5 and passes through the origin.
Answer by Cromlix(4381)

lines that are parallel to one
another have the same slope = 1
Using the equation:
y - b = m(x - a)
a = 0, b = 0 and m = 1
y - 0 = 1(x - 0)
y = x
This is the equation.
Hope this helps.
:-)